OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2

OpenStudy (anonymous):

first equn * by 2 8x + 2y = 16 adding 9x = 18 x = 2

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today?

OpenStudy (anonymous):

plz click good answer if u r happy with the answer

OpenStudy (anonymous):

Can you answer the second question i asked plz?

OpenStudy (amistre64):

By cramers rule ... matrix stuff 4x + y = 8 x − 2y = 2 x = 2+16 / 1+8 = 18 / 9 =2